OpenStudy (anonymous):

dp/dt=kp(L-P), where L is the upper limit of P and k is a constant. Prove that L and k are positive constant of p=L/1+Ce^(-klt)

OpenStudy (anonymous):

\[\frac{ dp }{ dt }=kP(L-P) P=\frac{ L }{ 1+Ce ^{-klt} }\]

OpenStudy (anonymous):

thanks for ur help

OpenStudy (anonymous):

sadly thats only one of the question i find really hard

OpenStudy (anonymous):

ill post the other question some other times

OpenStudy (loser66):

@chris00 @oldrin.bataku @eliassaab @tkhunny @druminjosh

OpenStudy (anonymous):

Can you separate the variables and set up the antiderivative you need to solve?

OpenStudy (loser66):

Can you do something for her/him? I have no clue.

OpenStudy (anonymous):

the only way to prove is to solve the equation. its a separable ODE (the logistic equation)

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

I got this part and posted it but the asker and me couldn't see the link between the problem and the solution. The question is Proving and we Solve. Ha!!

OpenStudy (anonymous):

they have similar equations

OpenStudy (anonymous):

yes but when you integrate and solve you will see that L has to be positive